# Homological stratification and descent

### 04 Mar 2026 v2

@@ Introduction
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a family of geometric functors that jointly detect when an object of $\mathscr{T}$ is zero. Suppose that $\mathscr{S}_{i}$ is tt-stratified for all $i\in I$. If $\mathscr{T}$ satisfies the Nerves of Steel conjecture and has a weakly noetherian spectrum, then the following conditions are equivalent:
@@
For a rigidly-compactly generated tt-category $\mathscr{T}$, the following are equivalent:
@@
Here $E_{\mathscr{B}}$ is a pure-injective object in $\mathscr{T}$ associated to each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$; see 1.1. The equivalence of the first two conditions is completely analogous to the characterization of cohomological stratification and tt-stratification established in [BIK11b, Theorem 4.2] and [BHS23d, Theorem 4.1], respectively. The third characterization is new and involves a notion of homological cosupport $\operatorname{Cosupp}^{h}$ which we introduce and study in this paper, motivated by its tt-theoretic counterpart studied in [BCH+23b].
When establishing that a given category is tt-stratified by checking the conditions analogous to part (b) of the above theorem, the bulk of the work goes into establishing minimality. In fact, for cohomological stratification — where the Zariski spectrum is noetherian — stratification is equivalent to minimality. In sharp contrast, the homological minimality condition should — heuristically speaking — be detected in the residue fields of $\mathscr{T}$ and thereby be much easier to verify in practice. This is the underlying reason why the theory of homological stratification tends to be better behaved under descent than its cousins.
For instance, in favourable situations, the points of $\operatorname{Spc}^{h}(\mathscr{T}^{c})$ are detected not only by homological residue fields, but by tt-residue fields of $\mathscr{T}$, i.e., for each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T})$ there exists a geometric functor $\mathscr{T}\to\mathscr{F}$ to a tt-field which detects $\mathscr{B}$. If this is the case, we say that $\mathscr{T}$ admits enough tt-fields (1.12). As demonstrated in [BC21], examples abound. Under the assumption that $\mathscr{T}$ admits enough residue fields, homological stratification reduces to the homological local-to-global principle:
@@
Suppose that $\mathscr{T}$ admits enough tt-fields $(f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}})$. Then the following are equivalent:
@@
The last condition says that the homological local-to-global principle holds in $\mathscr{T}$ whenever its tt-residue fields cover the entire category. It is possible to substantially generalize this to a relative context. To this end, we isolate a very general descent condition: A collection of geometric functors $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ is said to be weakly descendable (7.2) if the essential images of the corresponding right adjoints $(f_{i})_{*}$ generate $\mathscr{T}$ as a localizing ideal:
@@
Note that we do not require any smallness conditions on $(f_{i})_{*}(\mathbb{1}_{\mathscr{S}_{i}})\in\mathscr{T}$. This notion of weak descendability covers all examples of descent previously considered in the tt-literature, and is arguably the most general setting in which one could expect descent to hold (7.13). With this in mind, the following result establishes that h-stratification essentially always satisfies descent.
@@
This is in marked contrast to tt-stratification, where one has so far only been able to prove a patchwork of different descent theorems. With a satisfactory descent theorem for homological stratification in hand, one may then ask for a comparison with tt-stratification:
@@
If $\mathscr{T}$ is a rigidly-compactly generated tt-category with $\operatorname{Spc}(\mathscr{T}^{c})$ weakly noetherian, then the following are equivalent:
@@
Thus, whenever the Nerves of Steel conjecture is known or can be established through different methods, it suffices to consider the theory of homological stratification. This enables us to deduce strong novel descent results for tt-stratification itself. For instance, we prove:
@@
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable geometric family with each $\mathscr{S}_{i}$ tt-stratified. Assume additionally that $I$ is finite, that each right adjoint $(f_{i})_{*}$ preserves compact objects, and that each $\operatorname{Spc}(\mathscr{S}_{i}^{c})$ is noetherian. Then $\mathscr{T}$ is tt-stratified.
@@
Let $\mathscr{C}$ be a rigidly-compactly generated symmetric monoidal stable $\infty$-category and let $A\in\operatorname{CAlg}(\mathscr{C}^{c})$ be a descendable dualizable commutative algebra in $\mathscr{C}$. If $\operatorname{Mod}_{\mathscr{C}}(A)$ is tt-stratified and has a noetherian spectrum, then $\mathscr{C}$ is also tt-stratified and also has a noetherian spectrum.
@@
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable family of geometric functors with $\mathscr{S}_{i}$ tt-stratified for all $i\in I$ and let
@@
be the induced map on tt-spectra. Assume that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian. Then:
@@
    (Nil-descent) If $\varphi$ is injective then $\mathscr{T}$ is tt-stratified.
@@
    (Quasi-finite descent) If each right adjoint $(f_{i})_{*}$ preserves compact objects and $\varphi$ has discrete fibers then $\mathscr{T}$ is tt-stratified.
@@
In Section 1, we recall the definition and basic features of homological support and introduce the homological detection property. We also discuss the “naive” definition of homological support and establish situations in which it coincides with the homological support. Finally, we introduce homological cosupport and the homological codetection property. In Section 2, we introduce the homological local-to-global principle and discuss its relationship with the h-codetection property. This leads to the definition of h-stratification in Section 3 and the fundamental characterization of h-stratification in terms of the homological local-to-global principle and homological minimality. We discuss examples and counterexamples arising in commutative algebra and explain that h-stratification is equivalent to the homological local-to-global principle when there are enough tt-fields (3.9). We also show that a classification of Bousfield classes is implied by (and almost equivalent to) the h-detection property (3.18). This surprising result has no tt-analogue. Finally, in Section 4, we provide a further characterization of h-stratification in terms of h-cosupport (4.5).
We compare the homological support with the Balmer–Favi support in Section 5 and establish several base change results for the homological support in Section 6. Of particular note is 6.4, which establishes that the homological support satisfies the Avrunin–Scott identity whenever the h-detection property holds. In Section 7, we introduce the notion of a weakly descendable family of functors and argue that this is the appropriate context in which to study descent. This culminates in 7.22. We then turn to the comparison between h-stratification and tt-stratification in Section 8 yielding 8.6 and its corollaries. Then in Section 9 we turn to the task of descending the Nerves of Steel conjecture, yielding 9.2 and 9.4. We also show that the Nerves of Steel conjecture is equivalent to a purely support-theoretic statement (9.6). Our descent results are then applied in Section 10 to obtain several specific descent results for tt-stratification, including 10.3 among others. Applications to equivariant spectra are then given in Section 11, including 11.9. Finally, we conclude the paper in Section 12 with a list of open problems.
@@
Throughout $\mathscr{T}$ and $\mathscr{S}$ will denote rigidly-compactly generated tt-categories and $f^{*}\colon\mathscr{T}\to\mathscr{S}$ will denote a geometric functor.
  • We write $\varphi\colon\operatorname{Spc}(\mathscr{S}^{c})\to\operatorname{Spc}(\mathscr{T}^{c})$ for the induced map on tt-spectra and write $\varphi^{h}\colon\operatorname{Spc}^{h}(\mathscr{S}^{c})\to\operatorname{Spc}^{h}(\mathscr{T}^{c})$ for the induced map on homological spectra.
@@ 1. Homological support and cosupport
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category.
@@
A rigidly-compactly generated tt-category $\mathscr{T}$ admits enough tt-fields if for each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ there exists a geometric functor $\mathscr{T}\to\mathscr{F}$ to a tt-field, in the sense of [BKS19, Definition 1.1], which maps the unique point in $\operatorname{Spc}^{h}(\mathscr{F}^{c})$ to $\mathscr{B}$. Note that $\operatorname{Spc}^{h}(\mathscr{F}^{c})$ consists of a single point by [BKS19, Theorem 5.17(a)].
@@
The derived category $\operatorname{D}(R)$ of a commutative ring $R$ admits enough tt-fields. Indeed, $\operatorname{Spc}^{h}(\operatorname{D}(R)^{c})\cong\operatorname{Spec}(R)$ and tt-fields are provided by the usual residue fields $\operatorname{D}(R)\to\operatorname{D}(\kappa(\mathfrak{p}))$. For the homological prime $\mathscr{B}\in\operatorname{Spc}^{h}(\operatorname{D}(R)^{c})$ corresponding to $\mathfrak{p}\in\operatorname{Spec}(R)$, we have $E_{\mathscr{B}}\simeq\kappa(\mathfrak{p})$, considered as a complex concentrated in degree 0; see [BC21, Corollary 3.3].
@@
Let $f^{*}\colon\mathscr{T}\to\mathscr{F}$ be a geometric functor to a tt-field $\mathscr{F}$. Let $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ be the image of the unique point in $\operatorname{Spc}^{h}(\mathscr{F}^{c})$. Then $\mathscr{B}\in\operatorname{Supp}^{n}(t)$ if and only if $\mathscr{B}\in\operatorname{Supp}^{h}(t)$.
@@
If $\mathsf{hom}(t,E_{\mathscr{B}})=0$, then $\mathsf{hom}(f^{*}t,f^{!}E_{\mathscr{B}})=f^{!}\mathsf{hom}(t,E_{\mathscr{B}})=0$. Since $\mathscr{F}$ is a tt-field, it follows that $f^{*}t=0$ by [BKS19, Theorem 5.21]. Hence $f_{*}(\mathbb{1})\otimes t=0$ by the projection formula. Therefore $E_{\mathscr{B}}\otimes t=0$ since $E_{\mathscr{B}}$ is a direct summand of $f_{*}(\mathbb{1})$. ∎
@@
Suppose $\mathscr{T}$ admits enough tt-fields (1.12). Then
@@ 2. The homological local-to-global principle
The h-LGP property of 2.1 implies the equivalent conditions of 2.2. However, a priori the latter conditions are weaker since we do not know whether $\operatorname{Supp}^{h}(t)=\operatorname{Supp}^{n}(t)$ in general. In particular, it is not immediate from the above whether the h-codetection property implies the h-detection property. This is in contrast to the situation in [BCH+23b]. However, we have the following conditional result:
@@ 3. Homological stratification
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category. The following conditions are equivalent:
@@
A rigidly-compactly generated tt-category $\mathscr{T}$ is said to be homologically stratified (or h-stratified, for short) if it satisfies the equivalent conditions of 3.1.
Remark 3.3.
@@
Any tt-field $\mathscr{F}$ is h-stratified. Indeed, first note that $\mathscr{F}$ admits a unique homological prime by [BKS19, Theorem 5.17(a)]. It then suffices to show that the only localizing ideals of $\mathscr{F}$ are the zero ideal and $\mathscr{F}$ itself, that is, $\mathbb{1}\in\operatorname{Locid}\langle t\rangle$ for any nonzero $t\in\mathscr{F}$. By the definition of tt-field, any nonzero object $t$ admits a nonzero rigid-compact object $c$ as a retract, hence $\operatorname{Locid}\langle c\rangle\subseteq\operatorname{Locid}\langle t\rangle$. Moreover, $\mathbb{1}\in\operatorname{Thickid}\langle c\rangle\subseteq\operatorname{Locid}\langle c\rangle$ since $\operatorname{Spc}(\mathscr{F}^{c})$ is a singleton by [BKS19, Proposition 5.15]. Therefore, $\mathbb{1}\in\operatorname{Locid}\langle t\rangle$.
@@
Let $f^{*}\colon\mathscr{T}\to\mathscr{F}$ be a geometric functor to a tt-field and let $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ be the unique point in the image of $\varphi^{h}$. Assume that $\mathscr{T}$ has no nontrivial tensor-nilpotent objects. Then $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle f_{*}(\mathbb{1})\rangle$ is a minimal localizing ideal.
@@
If $\mathscr{T}$ has no nontrivial tensor-nilpotent objects and admits enough tt-fields, then it has h-minimality at all homological primes.
@@
The corollary applies whenever $\mathscr{T}$ has the h-detection property and admits enough tt-fields.
@@
Note that if $f^{*}\colon\mathscr{T}\to\mathscr{F}$ is a geometric functor to a tt-field then $f^{*}(t)=0$ for any tensor-nilpotent object $t\in\mathscr{T}$. Hence, the argument in the proof of 3.5 shows that if $t\in\operatorname{Locid}\langle f_{*}(\mathbb{1})\rangle$ is tensor-nilpotent then $t^{\otimes 2}=0$. In other words, this localizing ideal can only contain tensor-nilpotent objects of order at most 2.
@@
If $\mathscr{T}$ admits enough tt-fields then the following are equivalent:
@@
Certainly $(b)$ implies $(a)$ by 3.1. Conversely, h-LGP implies the h-detection property so, since we have enough tt-fields by assumption, we automatically get h-minimality by 3.6. ∎
@@
The corollary applies to the derived category $\mathscr{T}=\operatorname{D}(R)$ of any commutative ring since it admits enough tt-fields (1.13). Therefore, $\operatorname{D}(R)$ is h-stratified if and only if $R\in\operatorname{Locid}\langle\kappa(\mathfrak{p})\mid\mathfrak{p}\in\operatorname{Spec}(R)\rangle$, in light of 2.5.
@@
Let $\mathscr{T}=\operatorname{SH}$ be the stable homotopy category. By [BAL20b, Corollary 5.10], which relies on the nilpotence theorem of Devinatz–Hopkins–Smith [DHS88], the canonical map $\operatorname{Spc}^{h}(\operatorname{SH}^{c})\to\operatorname{Spc}(\operatorname{SH}^{c})$ is a bijection; cf. 8.1 below. In particular, for a given prime number $p$, there is a homological prime $\mathscr{B}_{p,\infty}$ corresponding to the tt-prime of all finite $p$-torsion spectra. The associated weak ring identifies with the mod $p$ Eilenberg–Mac Lane spectrum $H{\mathbb{F}_{\hskip-0.81949ptp}}$, i.e., $E_{\mathscr{B}_{p,\infty}}=H{\mathbb{F}_{\hskip-0.81949ptp}}$; see [BC21, Corollary 3.6]. However, it follows from Ravenel’s work [RAV84] that there are strict inclusions
@@
where $I$ denotes the $p$-local Brown–Comenetz dual of the sphere spectrum. Indeed, on the one hand, we have $I\in\operatorname{Locid}\langle H{\mathbb{F}_{\hskip-0.81949ptp}}\rangle$ because the homotopy groups of $I$ are $p$-torsion and concentrated in non-negative degrees. On the other hand, we have $H{\mathbb{F}_{\hskip-0.81949ptp}}\not\in\operatorname{Locid}\langle I\rangle$ because $H{\mathbb{F}_{\hskip-0.81949ptp}}\otimes H{\mathbb{F}_{\hskip-0.81949ptp}}\neq 0$ while $H{\mathbb{F}_{\hskip-0.81949ptp}}\otimes I=0$.
@@
Let $A$ be the truncated polynomial ring over a field $k$ considered by [NEE00]. We grade $A$ as in [DP08]. The derived category $\operatorname{D}(A)$ admits a unique tt-field $\operatorname{D}(A)\to\operatorname{D}(k)$, induced by the quotient $A\to k$ by the unique prime ideal of $A$. The pure-injective object $E_{\mathscr{B}}$ associated to the unique homological prime $\mathscr{B}$ is just the ordinary residue field $k$. Let $I=\operatorname{Hom}_{k}^{*}(A,k)$ be the graded $k$-dual of $A$, concentrated in non-positive degrees. We claim that $\operatorname{Locid}\langle I\rangle$ is strictly contained in $\operatorname{Locid}\langle k\rangle$. Indeed, we have $I\in\operatorname{Locid}\langle k\rangle$ by [DP08, Lemma 4.9]. On the other hand, $k\not\in\operatorname{Locid}\langle I\rangle$ since $I\otimes k=0$ by [DP08, Proposition 4.11] while $k\otimes k\neq 0$.
@@
Recall from 3.9 that the homological local-to-global principle implies h-minimality in the presence of enough tt-fields. Hence the homological local-to-global principle fails in the previous two examples. Indeed, we have $\operatorname{Supp}^{h}(I)=\varnothing$} in both examples above.

### 18 Mar 2026 v3

@@ Introduction
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a family of geometric functors that jointly detect when an object of $\mathscr{T}$ is zero. Suppose that $\mathscr{S}_{i}$ is $t$-stratified for all $i\in I$. If $\mathscr{T}$ satisfies the Nerves of Steel conjecture and has a weakly noetherian spectrum, then the following conditions are equivalent:
@@
For a rigidly-compactly generated $t$-category $\mathscr{T}$, the following are equivalent:
@@
Here $E_{\mathscr{B}}$ is a pure-injective object in $\mathscr{T}$ associated to each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$; see $1.$ The equivalence of the first two conditions is completely analogous to the characterization of cohomological stratification and $t$-stratification established in [BIK11b, Theorem 4.2] and [BHS23d, Theorem 4.1], respectively. The third characterization is new and involves a notion of homological cosupport $\operatorname{Cosupp}^{h}$ which we introduce and study in this paper, motivated by its $t$-theoretic counterpart studied in [BCH+23b].
When establishing that a given category is $t$-stratified by checking the conditions analogous to part (b) of the above theorem, the bulk of the work goes into establishing minimality. In fact, for cohomological stratification — where the Zariski spectrum is noetherian — stratification is equivalent to minimality. In sharp contrast, the homological minimality condition should — heuristically speaking — be detected in the residue fields of $\mathscr{T}$ and thereby be much easier to verify in practice. This is the underlying reason why the theory of homological stratification tends to be better behaved under descent than its cousins.
For instance, in favourable situations, the points of $\operatorname{Spc}^{h}(\mathscr{T}^{c})$ are detected not only by homological residue fields, but by $t$-residue fields of $\mathscr{T}$, i.e., for each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T})$ there exists a geometric functor $\mathscr{T}\to\mathscr{F}$ to a $t$-field which detects $\mathscr{B}$. If this is the case, we say that $\mathscr{T}$ admits enough $t$-fields (1.12). As demonstrated in [BC21], examples abound. Under the assumption that $\mathscr{T}$ admits enough residue fields, homological stratification reduces to the homological local-to-global principle:
@@
Suppose that $\mathscr{T}$ admits enough $t$-fields $(f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}})$. Then the following are equivalent:
@@
The last condition says that the homological local-to-global principle holds in $\mathscr{T}$ whenever its $t$-residue fields cover the entire category. It is possible to substantially generalize this to a relative context. To this end, we isolate a very general descent condition: A collection of geometric functors $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ is said to be weakly descendable (7.2) if the essential images of the corresponding right adjoints $(f_{i})_{*}$ generate $\mathscr{T}$ as a localizing ideal:
@@
Note that we do not require any smallness conditions on $(f_{i})_{*}(\mathbb{1}_{\mathscr{S}_{i}})\in\mathscr{T}$. This notion of weak descendability covers all examples of descent previously considered in the $t$-literature, and is arguably the most general setting in which one could expect descent to hold (7.13). With this in mind, the following result establishes that h-stratification essentially always satisfies descent.
@@
This is in marked contrast to $t$-stratification, where one has so far only been able to prove a patchwork of different descent theorems. With a satisfactory descent theorem for homological stratification in hand, one may then ask for a comparison with $t$-stratification:
@@
If $\mathscr{T}$ is a rigidly-compactly generated $t$-category with $\operatorname{Spc}(\mathscr{T}^{c})$ weakly noetherian, then the following are equivalent:
@@
Thus, whenever the Nerves of Steel conjecture is known or can be established through different methods, it suffices to consider the theory of homological stratification. This enables us to deduce strong novel descent results for $t$-stratification itself. For instance, we prove:
@@
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable geometric family with each $\mathscr{S}_{i}$ $t$-stratified. Assume additionally that II is finite, that each right adjoint $(f_{i})_{*}$ preserves compact objects, and that each $\operatorname{Spc}(\mathscr{S}_{i}^{c})$ is noetherian. Then $\mathscr{T}$ is $t$-stratified.
@@
Let $\mathscr{C}$ be a rigidly-compactly generated symmetric monoidal stable $\infty$-category and let $A\in\operatorname{CAlg}(\mathscr{C}^{c})$ be a descendable dualizable commutative algebra in $\mathscr{C}$. If $\operatorname{Mod}_{\mathscr{C}}(A)$ is $t$-stratified and has a noetherian spectrum, then $\mathscr{C}$ is also $t$-stratified and also has a noetherian spectrum.
@@
Let $(f_{i}^{*}\colon\mathscr{T}\to\mathscr{S}_{i})_{i\in I}$ be a weakly descendable family of geometric functors with $\mathscr{S}_{i}$ $t$-stratified for all $i\in I$ and let
@@
be the induced map on $t$-spectra. Assume that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian. Then:
@@
    (Nil-descent) If $\varphi$ is injective then $\mathscr{T}$ is $t$-stratified.
@@
    (Quasi-finite descent) If each right adjoint $(f_{i})_{*}$ preserves compact objects and $\varphi$ has discrete fibers then $\mathscr{T}$ is $t$-stratified.
@@
In Section 1, we recall the definition and basic features of homological support and introduce the homological detection property. We also discuss the “naive” definition of homological support and establish situations in which it coincides with the homological support. Finally, we introduce homological cosupport and the homological codetection property. In Section 2, we introduce the homological local-to-global principle and discuss its relationship with the h-codetection property. This leads to the definition of h-stratification in Section 3 and the fundamental characterization of h-stratification in terms of the homological local-to-global principle and homological minimality. We discuss examples and counterexamples arising in commutative algebra and explain that h-stratification is equivalent to the homological local-to-global principle when there are enough $t$-fields (3.9). We also show that a classification of Bousfield classes is implied by (and almost equivalent to) the h-detection property (3.18). This surprising result has no $t$-analogue. Finally, in Section 4, we provide a further characterization of h-stratification in terms of h-cosupport (4.5).
We compare the homological support with the Balmer–Favi support in Section 5 and establish several base change results for the homological support in Section 6. Of particular note is 6.4, which establishes that the homological support satisfies the Avrunin–Scott identity whenever the h-detection property holds. In Section 7, we introduce the notion of a weakly descendable family of functors and argue that this is the appropriate context in which to study descent. This culminates in 7.22. We then turn to the comparison between h-stratification and $t$-stratification in Section 8 yielding 8.6 and its corollaries. Then in Section 9 we turn to the task of descending the Nerves of Steel conjecture, yielding 9.2 and 9.4. We also show that the Nerves of Steel conjecture is equivalent to a purely support-theoretic statement (9.6). Our descent results are then applied in Section 10 to obtain several specific descent results for $t$-stratification, including 10.3 among others. Applications to equivariant spectra are then given in Section 11, including 11.9. Finally, we conclude the paper in Section 12 with a list of open problems.
@@
Throughout $\mathscr{T}$ and $\mathscr{S}$ will denote rigidly-compactly generated $t$-categories and $f^{*}\colon\mathscr{T}\to\mathscr{S}$ will denote a geometric functor.
  • We write $\varphi\colon\operatorname{Spc}(\mathscr{S}^{c})\to\operatorname{Spc}(\mathscr{T}^{c})$ for the induced map on $t$-spectra and write $\varphi^{h}\colon\operatorname{Spc}^{h}(\mathscr{S}^{c})\to\operatorname{Spc}^{h}(\mathscr{T}^{c})$ for the induced map on homological spectra.
@@ 1. Homological support and cosupport
Let $\mathscr{T}$ be a rigidly-compactly generated $t$-category.
@@
A rigidly-compactly generated $t$-category $\mathscr{T}$ admits enough $t$-fields if for each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ there exists a geometric functor $\mathscr{T}\to\mathscr{F}$ to a $t$-field, in the sense of [BKS19, Definition 1.1], which maps the unique point in $\operatorname{Spc}^{h}(\mathscr{F}^{c})$ to $\mathscr{B}$. Note that $\operatorname{Spc}^{h}(\mathscr{F}^{c})$ consists of a single point by [BKS19, Theorem 5.17(a)].
@@
The derived category $\operatorname{D}(R)$ of a commutative ring $R$ admits enough $t$-fields. Indeed, $\operatorname{Spc}^{h}(\operatorname{D}(R)^{c})\cong\operatorname{Spec}(R)$ and $t$-fields are provided by the usual residue fields $\operatorname{D}(R)\to\operatorname{D}(\kappa(\mathfrak{p}))$. For the homological prime $\mathscr{B}\in\operatorname{Spc}^{h}(\operatorname{D}(R)^{c})$ corresponding to $\mathfrak{p}\in\operatorname{Spec}(R)$, we have $E_{\mathscr{B}}\simeq\kappa(\mathfrak{p})$, considered as a complex concentrated in degree 0; see [BC21, Corollary 3.3].
@@
Let $f^{*}\colon\mathscr{T}\to\mathscr{F}$ be a geometric functor to a $t$-field $\mathscr{F}$. Let $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ be the image of the unique point in $\operatorname{Spc}^{h}(\mathscr{F}^{c})$. Then $\mathscr{B}\in\operatorname{Supp}^{n}(t)$ if and only if $\mathscr{B}\in\operatorname{Supp}^{h}(t)$.
@@
If $\mathsf{hom}(t,E_{\mathscr{B}})=0$, then $\mathsf{hom}(f^{*}t,f^{!}E_{\mathscr{B}})=f^{!}\mathsf{hom}(t,E_{\mathscr{B}})=0$. Since $\mathscr{F}$ is a $t$-field, it follows that $f^{*}t=0$ by [BKS19, Theorem 5.21]. Hence $f_{*}(\mathbb{1})\otimes t=0$ by the projection formula. Therefore $E_{\mathscr{B}}\otimes t=0$ since $E_{\mathscr{B}}$ is a direct summand of $f_{*}(\mathbb{1})$. ∎
@@
Suppose $\mathscr{T}$ admits enough $t$-fields (1.12). Then
@@ 2. The homological local-to-global principle
The h-LGP property of 2.1 implies the equivalent conditions of $2.$ However, a priori the latter conditions are weaker since we do not know whether $\operatorname{Supp}^{h}(t)=\operatorname{Supp}^{n}(t)$ in general. In particular, it is not immediate from the above whether the h-codetection property implies the h-detection property. This is in contrast to the situation in [BCH+23b]. However, we have the following conditional result:
@@ 3. Homological stratification
Let $\mathscr{T}$ be a rigidly-compactly generated $t$-category. The following conditions are equivalent:
@@
A rigidly-compactly generated $t$-category $\mathscr{T}$ is said to be homologically stratified (or h-stratified, for short) if it satisfies the equivalent conditions of 3.1.
Remark $3.$
@@
Any $t$-field $\mathscr{F}$ is h-stratified. Indeed, first note that $\mathscr{F}$ admits a unique homological prime by [BKS19, Theorem 5.17(a)]. It then suffices to show that the only localizing ideals of $\mathscr{F}$ are the zero ideal and $\mathscr{F}$ itself, that is, $\mathbb{1}\in\operatorname{Locid}\langle t\rangle$ for any nonzero $t\in\mathscr{F}$. By the definition of $t$-field, any nonzero object $t$ admits a nonzero rigid-compact object $c$ as a retract, hence $\operatorname{Locid}\langle c\rangle\subseteq\operatorname{Locid}\langle t\rangle$. Moreover, $\mathbb{1}\in\operatorname{Thickid}\langle c\rangle\subseteq\operatorname{Locid}\langle c\rangle$ since $\operatorname{Spc}(\mathscr{F}^{c})$ is a singleton by [BKS19, Proposition 5.15]. Therefore, $\mathbb{1}\in\operatorname{Locid}\langle t\rangle$.
@@
Let $f^{*}\colon\mathscr{T}\to\mathscr{F}$ be a geometric functor to a $t$-field and let $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$ be the unique point in the image of $\varphi^{h}$. Assume that $\mathscr{T}$ has no nontrivial tensor-nilpotent objects. Then $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle f_{*}(\mathbb{1})\rangle$ is a minimal localizing ideal.
@@
If $\mathscr{T}$ has no nontrivial tensor-nilpotent objects and admits enough $t$-fields, then it has h-minimality at all homological primes.
@@
The corollary applies whenever $\mathscr{T}$ has the h-detection property and admits enough $t$-fields.
@@
Note that if $f^{*}\colon\mathscr{T}\to\mathscr{F}$ is a geometric functor to a $t$-field then $f^{*}(t)=0$ for any tensor-nilpotent object $t\in\mathscr{T}$. Hence, the argument in the proof of 3.5 shows that if $t\in\operatorname{Locid}\langle f_{*}(\mathbb{1})\rangle$ is tensor-nilpotent then $t^{\otimes 2}=0$. In other words, this localizing ideal can only contain tensor-nilpotent objects of order at most 2.
@@
If $\mathscr{T}$ admits enough $t$-fields then the following are equivalent:
@@
Certainly $(b)$ implies $(a)$ by 3.1. Conversely, h-LGP implies the h-detection property so, since we have enough $t$-fields by assumption, we automatically get h-minimality by 3.6. ∎
@@
The corollary applies to the derived category $\mathscr{T}=\operatorname{D}(R)$ of any commutative ring since it admits enough $t$-fields (1.13). Therefore, $\operatorname{D}(R)$ is h-stratified if and only if $R\in\operatorname{Locid}\langle\kappa(\mathfrak{p})\mid\mathfrak{p}\in\operatorname{Spec}(R)\rangle$, in light of 2.5.
@@
Let $\mathscr{T}=\operatorname{SH}$ be the stable homotopy category. By [BAL20b, Corollary 5.10], which relies on the nilpotence theorem of Devinatz–Hopkins–Smith [DHS88], the canonical map $\operatorname{Spc}^{h}(\operatorname{SH}^{c})\to\operatorname{Spc}(\operatorname{SH}^{c})$ is a bijection; cf. 8.1 below. In particular, for a given prime number $p$, there is a homological prime $\mathscr{B}_{p,\infty}$ corresponding to the $t$-prime of all finite $p$-torsion spectra. The associated weak ring identifies with the mod $p$ Eilenberg–Mac Lane spectrum $H{\mathbb{F}_{\hskip-0.81949ptp}}$, i.e., $E_{\mathscr{B}_{p,\infty}}=H{\mathbb{F}_{\hskip-0.81949ptp}}$; see [BC21, Corollary 3.6]. However, it follows from Ravenel’s work [RAV84] that there are strict inclusions
@@
where II denotes the $p$-local Brown–Comenetz dual of the sphere spectrum. Indeed, on the one hand, we have $I\in\operatorname{Locid}\langle H{\mathbb{F}_{\hskip-0.81949ptp}}\rangle$ because the homotopy groups of II are $p$-torsion and concentrated in non-negative degrees. On the other hand, we have $H{\mathbb{F}_{\hskip-0.81949ptp}}\not\in\operatorname{Locid}\langle I\rangle$ because $H{\mathbb{F}_{\hskip-0.81949ptp}}\otimes H{\mathbb{F}_{\hskip-0.81949ptp}}\neq 0$ while $H{\mathbb{F}_{\hskip-0.81949ptp}}\otimes I=0$.
@@
Let AA be the truncated polynomial ring over a field $k$ considered by [NEE00]. We grade AA as in [DP08]. The derived category $\operatorname{D}(A)$ admits a unique $t$-field $\operatorname{D}(A)\to\operatorname{D}(k)$, induced by the quotient $A\to k$ by the unique prime ideal of AA. The pure-injective object $E_{\mathscr{B}}$ associated to the unique homological prime $\mathscr{B}$ is just the ordinary residue field $k$. Let $I=\operatorname{Hom}_{k}^{*}(A,k)$ be the graded $k$-dual of AA, concentrated in non-positive degrees. We claim that $\operatorname{Locid}\langle I\rangle$ is strictly contained in $\operatorname{Locid}\langle k\rangle$. Indeed, we have $I\in\operatorname{Locid}\langle k\rangle$ by [DP08, Lemma 4.9]. On the other hand, $k\not\in\operatorname{Locid}\langle I\rangle$ since $I\otimes k=0$ by [DP08, Proposition 4.11] while $k\otimes k\neq 0$.
@@
Recall from 3.9 that the homological local-to-global principle implies h-minimality in the presence of enough $t$-fields. Hence the homological local-to-global principle fails in the previous two examples. Indeed, we have $\operatorname{Supp}^{h}(I)=\varnothing$} in both examples above.